# Collapsing regular Riemannian foliations with flat leaves

### v1

@@ Acknowledgments
I thank Wilderich Tuschmann and Karsten Grove for discussions on $\mathsf{F}$-structures, and I thank Fernando Galaz-Garcída and Alexander Lytchak for discussions on collapse theory. I also thank John Harvey for comments that improved the proofs of the main theorems. Last I thank Jesús Núñez-Zimbrón and Jaime Santos for pointing me to the results about equivariant convergence.
@@ 2. Preliminaries
In this section we present the preliminaries necessary for our results. We begin by presenting the definition of a singular Riemannian foliation, then present the notion of the inifinitesimal foliation and the holonomy of a leaf. We then present the notions of $A$- and $B$-foliations. We also present the necessary results from the theories of convergence of Riemannian manifolds with curvature bounds, equivariant convergence of metric spaces, and metric spaces with a lower curvature bound.
@@ 2.1. Singular Riemannian foliations
For a closed singular Riemannian foliation $(M,g,\mathcal{F})$ and $\ell\leq\dim(\mathcal{F})$ we define
$\Sigma_{\ell}=\{p\in M\mid\dim(L_{p})=\ell\}.$
@@
for $\ell\in\{0,\ldots,\dim(\mathcal{F})\}$ the each connected component $C$ of $\Sigma_{\ell}$ is an embedded submanifold of $M$. We denote by $\Sigma^{p}$ the connected component of $\Sigma_{\ell}$ containing $p$, where $\ell=\dim(L_{p})$. The collection $\{\Sigma^{p}\mid p\in M\}$ give the canonical stratification of $M$ by $\mathcal{F}$. For the top dimension $\dim(\mathcal{F})$, the set $\Sigma_{\dim(\mathcal{F})}$ is open, dense and connected. We refer to it as the regular stratum, and denote it by $M_{\mathrm{reg}}=\Sigma_{\dim(\mathcal{F})}$. In the case when $M_{\mathrm{reg}}=M$ we say that the foliation $\mathcal{F}$ is regular. We also remark that, all other strata $\Sigma_{\ell}$ have codimension at least $2$ ([22, Section 2.3]).
@@
We equip the quotient space $M/\mathcal{F}$ with the quotient topology, making the quotient map $\pi\colon M\to M/\mathcal{F}$ continuous. We refer to $M/\mathcal{F}$ equipped with the quotient topology as the leaf space of $\mathcal{F}$. When the foliation is closed and $M$ is complete, the Riemannian distance on $M$ induces an inner metric, called the quotient metric on the leaf space, and it induces quotient topology. Moreover as pointed out in [26, p. 2943], for a closed foliation on a complete manifold, the leaf space equipped with the quotient metric is locally an Alexandrov space (see Section 2.5 for a definition of an Alexandrov space).
@@
Given a singular Riemannian foliation $\mathcal{F}$ on a vector space $V$ equipped with an inner product $g$, we say that $(V,\mathcal{F},g)$ is an infinitesimal foliation if $(V,\mathcal{F},g)$ is a singular Riemannian foliation, and for the origin $\bar{0}\in V$ we have $L_{\bar{0}}=\{\bar{0}\}$.
@@
We now consider $(M,\mathcal{F},g)$ a singular Riemannian foliation on a complete Riemannian manifold, and fix $p\in M$. We denote by $\nu_{p}(M,L_{p})==\{v\in T_{p}M\mid g(p)(v,x)=0\mbox{ for any }x\in T_{p}L_{p}\}$ the set of normal space to the leaf at $p$. Given $\varepsilon>0$ we consider the closed normal disk of radius $\varepsilon$ $\nu_{p}^{\varepsilon}(M,L_{p})=\{v\in\nu_{p}(M,L_{p})\mid\|v\|^{2}_{g}\leq\varepsilon\}$. Then by fixing $\varepsilon>0$ small enough (for example $\varepsilon$ smaller than the injectivity radius at $p$), we may assume that $\operatorname{exp}_{p}\colon\nu_{p}^{\varepsilon}(M,L_{p})\to S_{p}=\exp_{p}(\nu_{p}^{\varepsilon}(M,L_{p}))\subset M$ is a diffeomorphism, where
$\nu_{p}^{\varepsilon}(M,L_{p})=\{v\in T_{p}M\mid g(p)(v,x)=0\mbox{ for any }x\in T_{p}L_{p},\>\|v\|^{2}\leq\varepsilon\}.$
For $v\in\nu_{p}^{\varepsilon}(M,L_{p})$ we define $\mathcal{L}_{v}\subset\nu_{p}^{\varepsilon}(M,L_{p})$ to be the connected component of $\operatorname{exp}_{p}^{-1}(L_{\operatorname{exp}_{p}(v)}\cap S_{p})$ containing $v$. In this way, we obtain a partition $\mathcal{F}_{p}(\varepsilon)$ on $\nu_{p}^{\varepsilon}(M,L_{p})$. By [35, Lemma 6.2] the foliation $\mathcal{F}_{p}(\varepsilon)$ is invariant under homotheties $h_{\lambda}\colon\nu_{p}^{\varepsilon}(M,L_{p})\to\nu_{p}^{\lambda\varepsilon}(M,L_{p})$, with $h_{\lambda}(v)=\lambda v$. From this we see that $\mathcal{F}(\varepsilon)$ is independent of $\varepsilon$, and thus we can extend $\mathcal{F}_{p}(\varepsilon)$ to a partition $\mathcal{F}_{p}$ on $\nu_{p}(M,L_{p})$ in a unique way. Moreover, [35, Proposition 6.5] states that $(\nu_{p}(M,L_{p}),\mathcal{F}_{p},g_{p}^{\perp})$ is an infinitesimal foliation, where $g_{p}^{\perp}=g(p)|_{\nu_{p}(M,L_{p})}$. We refer to $(\nu_{p}(M,L_{p}),\mathcal{F}_{p},g_{p}^{\perp})$ as the infinitesimal foliation of $\mathcal{F}$ at $p$.
@@
þ Let $(M,\mathcal{F},g)$ be a singular Riemannian foliation on an $m$-dimensional manifold, and for any $p\in M$ we write $k_{p}=\dim(L_{p})$. Then for $p\in M$ fixed there exists a coordinate system $(W,\phi)$, with $W\subset M$ and $\phi\colon W\to\mathbb{R}^{m}$ such that:
@@ 2.2. Holonomy
In this section we present for $(M,\mathcal{F},g)$ a singular Riemannian foliation the notion of the holonomy of a closed leaf $L_{p}\in\mathcal{F}$. We denote by $\nu(L)\to L$ the normal bundle to $L$, and we have the following construction:
@@
þ Let $L$ be a closed leaf of $(M,\mathcal{F},g)$ a singular Riemannian foliation, and let $\gamma\colon[0,1]\to L$ a piecewise smooth curve with $\gamma(0)=p$. Then there exists a map $G\colon[0,1]\times\nu_{p}(M,L)\to\nu(L)$ such that:
@@
Let $\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})$ denote the group of foliated isometries of the infinitesimal foliation. That is, a map $f\in\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})$ is an isometry $f\colon(\nu_{p}(M,L_{p}),g_{p}^{\perp})\to(\nu_{p}(M,L_{p}),g_{p}^{\perp})$ such that for any leaf $\mathcal{L}\in\mathcal{F}_{p}$ we have $f(\mathcal{L})\in\mathcal{F}_{p}$. We denote by $\mathrm{O}(\mathcal{F}_{p})\subset\mathrm{O}(\nu_{p}^{\perp}(M,L_{p}),\mathcal{F}_{p})$ the subgroup of foliated isometries that leave the infinitesimal foliation invariant, i.e. $f\in\mathrm{O}(\mathcal{F}_{p})$ if and only if $f(\mathcal{L})\subset\mathcal{L}$ for all $\mathcal{L}\in\mathcal{F}_{p}$. Observe that for a closed piecewise loop $\gamma\colon[0,1]\to L_{p}$ with start point $p$ we have by þ2.2 a map $G_{\gamma}\colon\nu_{p}(M,L_{p})\to\nu_{p}(M,L_{p})$ with $G_{\gamma}\in\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})$. Moreover, from [8, Proposition 2.5] it follows that the coset $G_{\gamma}\mathrm{O}(\mathcal{F}_{p})\in\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})/\mathrm{O}(\mathcal{F}_{p})$ depends only on the homotopy class of $\gamma$. Thus, we have a well defined map $\rho\colon\pi_{1}(L_{p},p)\to\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})/\mathrm{O}(\mathcal{F}_{p})$.
@@ 2.3. $A$- and $B$-foliations
þ Let $(M,\mathcal{F},g)$ be an $A$-foliation on a compact simply-connected manifold, and let $L\in\mathcal{F}$ be a regular leaf. Then $L$ is homeomorphic to a torus.
@@
We recall that for $n\geq 5$ by [31], there exists $n$-manifolds which are homoeomorphic to the $n$-torus, but not diffeomorphic. These manifolds are known as exotic tori. We recall that there are examples of $A$-foliations $(M,\mathcal{F},g)$ on compact manifolds whose leaves are exotic tori, but $M$ is not simply connected (see [16]).
In [22] the authors also introduced the concept of $B$-foliations, which are $A$-foliations whose leaves are homeomorphic to Bieberbach manifolds. Recall that an $n$-dimensional Bieberbach manifold $L$ is an aspherical manifold such that $\pi_{1}(L)$ is isomorphic to a discrete subgroup $G$ of $\mathbb{R}^{n}\rtimes\mathrm{O}(n)\subset\mathrm{Aff}(n)$ which is torsion free, and such that $\mathbb{R}^{n}/G$ is compact. Such a group $G$ is called a Bieberbach group. We recall that Bieberbach manifolds are those homeomorphic to smooth manifolds that admit a flat Riemannian metric, i.e. a Riemannian metric whose sectional curvature is equal to 00 for any $2$-plane in the tangent space at any point. By the following result, it follows that the leaves of $A$-foliations on compact simply-connected manifolds have fundamental group isomorphic to a Bieberbach group, and except for leaves of dimension $4$, they are homeomorphic to Bieberbach manifolds.
@@
This implies that, except for the leaves of dimension $4$, from a topological viewpoint $A$-foliations and the $B$-foliations as defined in [22] are indistinguishable on compact simply-connected manifolds. Thus in [8, Remark 3.4] the author proposed to define $B$-foliations as those $A$-foliations $(M,\mathcal{F},g)$ such that for any $L\in\mathcal{F}$ we have that $(L,g|_{L})$ is flat. This definition immediately introduces some rigidity to the diffeomorpshism type of the foliation: Combining þ2.3 with the work of Bieberbach [3] (and more generally by [15, Theorem 3]) we obtain that for simply-connected manifolds the regular leaves of an $A$-foliation $(M,\mathcal{F},g)$ such that $(L,g|_{L})$ is flat, i.e. $B$-foliation as defined in this paragraph, are diffeomorphic to the standard torus.
@@ 2.4. Compact flat metric spaces
For the sake of completeness, we give a short review of the geometry of flat manifolds. We consider $(M,g)$ to be a compact smooth Rieamannian $m$-dimensional manifold, such that at every point $p\in M$ and any $2$-dimensional plane in $T_{p}M$ we have the sectional curvature equal to 00, i.e. a flat manifold. By the Cartan-Hadamard theorem (see [12, Theorem 3.1]) it follows that the universal cover $\widetilde{M}$ of $M$ is diffeomorphic to $\mathbb{R}^{m}$. We consider $\widetilde{g}=\pi^{\ast}(g)$ the pull back metric of $g$ under the covering map $\pi\colon\widetilde{M}\to M$. Observe that by construction $\pi\colon(\widetilde{M},\widetilde{g})\to(M,g)$ is a local isometry, and thus $(\widetilde{M},\widetilde{g})$ is a flat manifold diffeomorphic to $\mathbb{R}^{m}$. By [12, Theorem 4.1] it follows that $(\widetilde{M},\widetilde{g})$ is isometric to $(\mathbb{R}^{m},\sigma)$ where $\sigma$ denotes the Euclidean Riemannian metric. Thus we have that there exists a covering map $\bar{\pi}\colon(\mathbb{R}^{m},\sigma)\to(M,g)$ which is a local isometry, i.e. $\bar{\pi}^{\ast}(g)=\sigma$. By letting $B$ be the group of deck transformations of $\bar{\pi}$, it follows that $B$ is a subgroup of the isometry group $\operatorname{Isom}(\mathbb{R}^{m},\sigma)$ which is torsion free, i.e. $B$ is a Bieberbach group. Thus $(M,g)$ is isometric to the homogeneous Riemannian space $(\mathbb{R}^{m}/B,\bar{\sigma})$.

### v2

@@ Acknowledgments
I thank Wilderich Tuschmann and Karsten Grove for discussions on $\mathsf{F}$-structures, and I thank Fernando Galaz-Garcída and Alexander Lytchak for discussions on collapse theory. I also thank John Harvey for comments that improved the proofs of the main theorems. Last I thank Jesús Núñez-Zimbrón and Jaime Santos for pointing me to the results about equivariant convergence. I also thank the anonymous referee for suggestions that improved the presentation of the proofs.
@@ 2. Preliminaries
In this section we present the preliminaries necessary for our results. We begin by presenting the definition of a singular Riemannian foliation, then present the notion of the infinitesimal foliation and the holonomy of a leaf. We then present the notions of $A$- and $B$-foliations. We also present the necessary results from the theories of convergence of Riemannian manifolds with curvature bounds, equivariant convergence of metric spaces, and metric spaces with a lower curvature bound.
@@ 2.1. Singular Riemannian foliations
For a closed singular Riemannian foliation $(M,g,\mathcal{F})$ and $0\leq\ell\leq\dim(\mathcal{F})$ we define the $\ell$-dimensional stratum as
$\Sigma_{\ell}=\{p\in M\mid\dim(L_{p})=\ell\},$
@@
and point out that each connected component $C$ of $\Sigma_{\ell}$ is an embedded submanifold of $M$ (see [45]). We denote by $\Sigma^{p}$ the connected component of $\Sigma_{\ell}$ containing $p$, where $\ell=\dim(L_{p})$. The collection $\{\Sigma^{p}\mid p\in M\}$ give the canonical stratification of $M$ by $\mathcal{F}$. For the top dimension $\dim(\mathcal{F})$, the set $\Sigma_{\dim(\mathcal{F})}$ is open, dense and connected. We refer to it as the regular stratum, and denote it by $M_{\mathrm{reg}}=\Sigma_{\dim(\mathcal{F})}$. In the case when $M_{\mathrm{reg}}=M$ we say that the foliation $\mathcal{F}$ is regular. We also remark that all other strata $\Sigma_{\ell}$ have codimension at least $2$ ([22, Section 2.3]).
@@
We equip the quotient space $M/\mathcal{F}$ with the quotient topology, making the quotient map $\pi\colon M\to M/\mathcal{F}$ continuous. We refer to $M/\mathcal{F}$ equipped with the quotient topology as the leaf space of $\mathcal{F}$. When the foliation is closed and $M$ is complete, the Riemannian distance on $M$ induces an inner metric, called the quotient metric on the leaf space, and it induces the quotient topology. Moreover as pointed out in [27, p. 2943], for a closed foliation on a complete manifold, the leaf space equipped with the quotient metric is locally an Alexandrov space (see Section 2.5 for a definition of an Alexandrov space).
@@
Given a singular Riemannian foliation $\mathcal{F}$ on a vector space $V$ equipped with an inner product $g$, we say that $(V,\mathcal{F},g)$ is an infinitesimal foliation if for the origin $\bar{0}\in V$ we have $L_{\bar{0}}=\{\bar{0}\}$.
@@
We now consider $(M,\mathcal{F},g)$ a singular Riemannian foliation on a complete Riemannian manifold, and fix $p\in M$. We denote by $\nu_{p}(M,L_{p})=\{v\in T_{p}M\mid g(p)(v,x)=0\mbox{ for any }x\in T_{p}L_{p}\}$ the set of normal space to the leaf at $p$. Given $\varepsilon>0$ we consider the closed normal disk of radius $\varepsilon$ defined as $\nu_{p}^{\varepsilon}(M,L_{p})=\{v\in\nu_{p}(M,L_{p})\mid\|v\|^{2}_{g}\leq\varepsilon\}$. Then by fixing $\varepsilon>0$ small enough (for example $\varepsilon$ smaller than the injectivity radius at $p$), we may assume that $\operatorname{exp}_{p}\colon\nu_{p}^{\varepsilon}(M,L_{p})\to S_{p}=\exp_{p}(\nu_{p}^{\varepsilon}(M,L_{p}))\subset M$ is a diffeomorphism. For $v\in\nu_{p}^{\varepsilon}(M,L_{p})$ we define $\mathcal{L}_{v}\subset\nu_{p}^{\varepsilon}(M,L_{p})$ to be the connected component of $\operatorname{exp}_{p}^{-1}(L_{\operatorname{exp}_{p}(v)}\cap S_{p})$ containing $v$. In this way, we obtain a partition $\mathcal{F}_{p}(\varepsilon)$ on $\nu_{p}^{\varepsilon}(M,L_{p})$. By [38, Lemma 6.2] the foliation $\mathcal{F}_{p}(\varepsilon)$ is invariant under homotheties $h_{\lambda}\colon\nu_{p}^{\varepsilon}(M,L_{p})\to\nu_{p}^{\lambda\varepsilon}(M,L_{p})$, with $h_{\lambda}(v)=\lambda v$. From this we see that $\mathcal{F}(\varepsilon)$ is independent of $\varepsilon$, and thus we can extend $\mathcal{F}_{p}(\varepsilon)$ to a partition $\mathcal{F}_{p}$ on $\nu_{p}(M,L_{p})$ in a unique way. Moreover, [38, Proposition 6.5] states that $(\nu_{p}(M,L_{p}),\mathcal{F}_{p},g_{p}^{\perp})$ is an infinitesimal foliation, where $g_{p}^{\perp}=g(p)|_{\nu_{p}(M,L_{p})}$. We refer to $(\nu_{p}(M,L_{p}),\mathcal{F}_{p},g_{p}^{\perp})$ as the infinitesimal foliation of $\mathcal{F}$ at $p$.
@@
Let $(M,\mathcal{F},g)$ be a singular Riemannian foliation on an $m$-dimensional manifold, and for any $p\in M$ we write $k_{p}=\dim(L_{p})$. Then for $p\in M$ fixed there exists a coordinate system $(W,\phi)$, with $W\subset M$ and $\phi\colon W\to\mathbb{R}^{m}$ such that:
@@ 2.2. Holonomy
In this section we present for $(M,\mathcal{F},g)$ a singular Riemannian foliation the notion of the holonomy of a closed leaf $L_{p}\in\mathcal{F}$ which we simply denote by $L$. We denote by $\nu(L)\to L$ the normal bundle to $L$, and we have the following construction:
@@
Let $L$ be a closed leaf of $(M,\mathcal{F},g)$ a singular Riemannian foliation, and let $\gamma\colon[0,1]\to L$ a piecewise smooth curve with $\gamma(0)=p$. Then there exists a map $G\colon[0,1]\times\nu_{p}(M,L)\to\nu(L)$ such that:
@@
Let $\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})$ denote the group of foliated isometries of the infinitesimal foliation. That is, a map $f\in\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})$ is an isometry $f\colon(\nu_{p}(M,L_{p}),g_{p}^{\perp})\to(\nu_{p}(M,L_{p}),g_{p}^{\perp})$ such that for any leaf $\mathcal{L}\in\mathcal{F}_{p}$ we have $f(\mathcal{L})\in\mathcal{F}_{p}$. We denote by $\mathrm{O}(\mathcal{F}_{p})\subset\mathrm{O}(\nu_{p}^{\perp}(M,L_{p}),\mathcal{F}_{p})$ the subgroup of foliated isometries that leave the infinitesimal foliation invariant, i.e. $f\in\mathrm{O}(\mathcal{F}_{p})$ if and only if $f(\mathcal{L})\subset\mathcal{L}$ for all $\mathcal{L}\in\mathcal{F}_{p}$. Observe that for a closed piecewise loop $\gamma\colon[0,1]\to L_{p}$ with start point $p$ we have by ?THM? LABEL:T:_holonomy_map_of_curve a map $G_{\gamma}\colon\nu_{p}(M,L_{p})\to\nu_{p}(M,L_{p})$ with $G_{\gamma}\in\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})$. Moreover, from [8, Proposition 2.5] it follows that the coset $G_{\gamma}\mathrm{O}(\mathcal{F}_{p})\in\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})/\mathrm{O}(\mathcal{F}_{p})$ depends only on the homotopy class of $\gamma$. Thus, we have a well defined map $\rho\colon\pi_{1}(L_{p},p)\to\mathrm{O}(\nu_{p}(M,L_{p}),\mathcal{F}_{p})/\mathrm{O}(\mathcal{F}_{p})$.
@@ 2.3. $A$- and $B$-foliations
Let $(M,\mathcal{F},g)$ be an $A$-foliation on a compact simply-connected manifold, and let $L\in\mathcal{F}$ be a regular leaf. Then $L$ is homeomorphic to a torus.
@@
We recall that for $n\geq 5$ by [33], there exists $n$-manifolds which are homeomorphic to the $n$-torus, but not diffeomorphic. These manifolds are known as exotic tori. We recall that there are examples of $A$-foliations $(M,\mathcal{F},g)$ on compact manifolds whose leaves are exotic tori, but $M$ is not simply connected (see [16]).
In [22] the authors also introduced the concept of $B$-foliations, which are $A$-foliations whose leaves are homeomorphic to Bieberbach manifolds. Recall that an $n$-dimensional Bieberbach manifold $L$ is an aspherical manifold such that $\pi_{1}(L)$ is isomorphic to a discrete subgroup $G$ of $\mathbb{R}^{n}\rtimes\mathrm{O}(n)\subset\mathrm{Aff}(n)$ which is torsion free, and such that $\mathbb{R}^{n}/G$ is compact. Such a group $G$ is called a Bieberbach group. We recall that Bieberbach manifolds are those homeomorphic to smooth manifolds that admit a flat Riemannian metric, i.e. a Riemannian metric whose sectional curvature is equal to 0 for any $2$-plane in the tangent space at any point. By the following result, we conclude that the leaves of $A$-foliations on compact simply-connected manifolds have fundamental group isomorphic to a Bieberbach group, and except for leaves of dimension $4$, they are homeomorphic to Bieberbach manifolds.
@@
This implies that, except for the leaves of dimension $4$, from a topological viewpoint $A$-foliations and the $B$-foliations as defined in [22] are indistinguishable on compact simply-connected manifolds. Thus in [8, Remark 3.4] the author proposed to define $B$-foliations as those $A$-foliations $(M,\mathcal{F},g)$ such that for any $L\in\mathcal{F}$ we have that $(L,g|_{L})$ is flat. This definition immediately introduces some rigidity to the diffeomorphism type of the foliation: Combining ?THM? LABEL:T:_regular_leaves_are_tori with the work of Bieberbach [3] (and more generally by [15, Theorem 3]) we obtain that for simply-connected manifolds the regular leaves of an $A$-foliation $(M,\mathcal{F},g)$ such that $(L,g|_{L})$ is flat, i.e. $B$-foliation as defined in this paragraph, are diffeomorphic to the standard torus.
Remark 2.6.
We point out that a singular Riemannian foliation $(M,\mathcal{F},g)$ induced by an effective smooth torus action $\mu\colon T^{k}\times M\to M$ by isometries is a $B$-foliation, in the sense that the orbits with the induced Riemannian metric are flat. We sketch briefly the proof of this fact. Fix $p\in M$ such that $T^{k}(p)$ is a principal orbit. Then for the isotropy group at $p$ we have $T^{k}_{p}=\{e\}$, where $e\in T^{k}$ is the identity element (see [48, Example 3.4.3]). This implies that the orbit $T^{k}(p)$ is equivariantly diffeomorphic to $T^{k}$ via the map $\mu_{p}\colon T^{k}\to T^{k}(p)$ given by $\mu_{p}(\xi)=\mu(\xi,p)$. Via this diffeomorphism the $T^{k}$-invariant induced metric $g|_{T^{k}(p)}$ induces a left-invariant metric $\bar{g}_{p}$ on $T^{k}$. Since $T^{k}$ is abelian, by [36, Corollary 1.3] we have that the sectional curvature of $\bar{g}_{p}$ is non-negative. If it is positive at some point (and thus positive anywhere), then by the Bonnet-Myers Theorem ([12, Chapter 9, 3.1 Theorem]) we have that $T^{k}$ has finite fundamental group, which is a contradiction. Thus, $\bar{g}_{p}$ is flat, and consequently $(T^{k}(p),g|_{T^{k}(p)})$ is a flat manifold. Consider now $q\in M$ not in a principal orbit. We point out that a connected component $(T^{k}_{q})^{0}$ of the identity element of the isotropy subgroup at $q$ is homeomorphic to a subtorus of $T^{k}$. Then there exists a finite cover $\overline{T^{k}(q)}$ of $T^{k}(q)$ which is homeomorphic to a torus of dimension equal to $k-\dim((T^{k}_{q})^{0})$ (see [22, Example 2.4]). The metric $g|_{T^{k}(q)}$ lifts to a $T^{k}/(T^{k}_{q})^{0}$-invariant metric $\tilde{g}_{q}$ on $\overline{T^{k}(q)}$. Observing that the metric $\tilde{g}_{q}$ induces a left-invariant metric on the torus $T^{k}/(T^{k}_{q})^{0}$, then by applying the same reasoning as for the principal orbit case we conclude that the metric $\tilde{g}_{q}$ is flat. Since $\overline{T^{k}(q)}$ is a finite cover we conclude that $g|_{T^{k}(q)}$ is flat.
@@ 2.4. Compact flat metric spaces
[…]